# Testing Noise Assumptions of Learning Algorithms

### v1

@@ Abstract
To model this question, we extend the recently proposed testable learning framework of Rubinfeld and Vasilyan [RV23] and require a learner to run an associated test that satisfies the following two conditions: (1) whenever the test accepts, the learner outputs a classifier along with a certificate of optimality, and (2) the test must pass for any dataset drawn according to a specified modeling assumption on both the marginal distribution and the noise model. We then consider the problem of learning halfspaces over Gaussian marginals with Massart noise (where each label can be flipped with probability less than $1/2$ depending on the input features), and give a fully-polynomial time testable learning algorithm.
@@ 1 Introduction
Developing efficient algorithms for learning in the presence of noise is one of the most fundamental problems in machine learning with a long line of celebrated research. Assumptions on the noise model itself vary greatly. For example, the well-studied random classification noise model (RCN) assumes that the label corruption process is independent across examples, whereas malicious noise models allow a fraction of the (joint) data-generating distribution to be changed adversarially. Understanding the computational landscape of learning with respect to different noise models remains a challenging open problem.
@@
We use the recently introduced testable learning [RV23] framework to model these questions. In this framework, a learner first runs a test on the training set. Whenever the test accepts, the learner outputs a classifier along with a proof that the classifier has near-optimal error. Furthermore, the test must accept with high probability whenever the training set is drawn from a distribution satisfying some specified set of modeling assumptions. If the test rejects, the learner recognizes that one of the modeling assumptions has failed and will therefore refrain from outputting a classifier. Here, our modeling assumptions will include both the structure of the noise model and the structure of the marginal distribution from which the data is generated.
More concretely, we will consider the problem of learning halfspaces under Gaussian marginals with respect to Massart noise, an extensively studied problem where an adversary flips binary labels independently with probability at most $1/2$ (the probability of flipping can vary across instances). The goal is to find a halfspace $\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})$ with near-optimal misclassification error rate $\mathsf{opt}+\epsilon$, where $\mathsf{opt}$ is the best misclassification error rate achievable by a halfspace. For this problem, a long line of work [ABHU15, ABHZ16, YZ17, ZLC17, MV19, DKTZ20a] resulted in the algorithm of Diakonikolas et al. [DKTZ20a] that runs in time $\mathrm{poly}(d/\epsilon)$ and achieves the optimal error rate. In contrast, the worst-case-noise version of this problem (i.e. agnostic learning) is believed to require exponential time in the accuracy parameter, even with respect to Gaussian marginals [KKMS08, DKK+21, GGK20].
@@ 1.1 Our Results
Let $f:\mathbb{R}^{d}\to\{\pm 1\}$ be a concept, let $\eta:\mathbb{R}^{d}\to[0,1/2]$ and let $\mathcal{D}$ be a distribution over $\mathbb{R}^{d}$. The oracle $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{D},f,\eta}$ receives $m\in\mathbb{N}$ and returns $m$ i.i.d. examples of the form $(\mathbf{x},y)\in\mathbb{R}^{d}\times\{\pm 1\}$, where $\mathbf{x}\sim\mathcal{D}$ and $y=\xi\cdot f(\mathbf{x})$, with $\xi=1$ w.p. $\eta(\mathbf{x})$ and $\xi=-1$ w.p. $1-\eta(\mathbf{x})$. The quantity $\sup_{\mathbf{x}\in\mathbb{R}^{d}}\eta(\mathbf{x})\in[0,1/2]$ is called the noise rate.
@@
Learning Setting. Our results work in the following extension of testable learning [RV23].
Definition 1.3 (Testable Learning, extension of Definition 4 in [RV23]).
@@
The difference between Definition 1.3 and the definition of [RV23] is that the completeness criterion does not only concern the marginal distribution on $\mathbb{R}^{d}$, but the joint distribution over $\mathbb{R}^{d}\times\{\pm 1\}$. The choice of the oracle class $\mathcal{O}$ encapsulates all of the modeling assumptions under which our algorithm should accept (both on the marginal distribution on $\mathbb{R}^{d}$, as well as on the labels). Note that the probability of success can be amplified through repetition (see [RV23]), so it suffices to solve the problem for $\delta=1/3$. Our main results and their relation to prior work are summarized in Table 1.
@@
Massart, $\eta_{0}=1/2-c$ $\mathrm{poly}(d,1/\epsilon)$ [DKTZ20a] $\mathrm{poly}(d,1/\epsilon)$ [Theorem 2.7]
@@
$d^{O(\log(1/\epsilon))}2^{\mathrm{poly}(1/\epsilon)}$ [DKK+22]
$d^{\Omega(\log(1/\epsilon))}$ [DKK+22]
$d^{\tilde{O}(1/\epsilon^{2})}$ [RV23, GKK23]
@@
$d^{\tilde{O}(1/\epsilon^{2})}$ [KKMS08]
$d^{\Omega(1/\epsilon^{2})}$ [DKPZ21]
$d^{\tilde{O}(1/\epsilon^{2})}$ [RV23, GKK23]
@@
Upper Bound. In Theorem 2.7, we show that there is a polynomial-time tester-learner for the class $\mathcal{H}_{\mathsf{hs}}$ with respect to $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},\mathcal{H}_{\mathsf{hs}},\eta_{0}}$ for any $\eta_{0}\leq 1/2-c$, where $c$ is any positive constant. Moreover, whenever our algorithm accepts, it is guaranteed to output the optimal halfspace with respect to the input dataset $\bar{S}$, even if $\bar{S}$ is not generated from i.i.d. examples and can, therefore, be completely arbitrary. Given the upper bounds of Table 1 our algorithm can be used as a first step before applying the more powerful (but, also, more expensive) tester-learner of [RV23, GKK23]. If our algorithm accepts, this might lead to significant, assumption-free computational savings.
Lower Bounds. Our upper bound holds when the noise rate is bounded away below $1/2$. We show that this is necessary: in the high-noise regime ($\eta_{0}=1/2$), the best known lower bounds for learning under adversarial label noise also hold in the testable setting, with respect to random classification noise of rate $1/2$ (Definition 2.1), which is a special version of Massart noise. We give both cryptographic lower bounds (Theorem 3.4) assuming subexponential hardness on the problem of learning with errors (LWE), as well as statistical query lower bounds (Theorem 3.6). Our lower bounds follow directly from lower bounds from the literature of agnostic learning [DKPZ21, Tie23, DKR23] (combined with the simple 3.1).
@@ 1.2 Our Techniques
An outline of the proof of Theorem 2.7 is provided in Section 2. The algorithm begins by running a halfspace learner that is guaranteed to output a near-optimal halfspace when the marginal is Gaussian and the noise is Massart. The main challenge is to certify that the candidate output $\mathbf{v}\in\mathbb{S}^{d-1}$ corresponds to a halfspace with near-optimal error. To this end, we carefully analyze the error of $\mathbf{v}$ (see Section 2.1) in terms of quantities that we can efficiently test using the following testers.
Disagreement tester and sandwiching polynomials. Suppose $\mathbf{v}$ and $\mathbf{v}^{\prime}$ are two unit vectors. The probability that $\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})$ equals $\measuredangle(\mathbf{v},\mathbf{v}^{\prime})/\pi$ when $\mathbf{x}$ is Gaussian. We show that this property can be approximately certified when only one of the two vectors, $\mathbf{v}$, is known. In particular, given a set $S$ of points in $\mathbb{R}^{d}$ and a vector $\mathbf{v}\in\mathbb{S}^{d-1}$, we can certify that for all $\mathbf{v}^{\prime}\in\mathbb{S}^{d-1}$ the fraction of points in $S$ such that $\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})$ is $(1\pm 0.1)\measuredangle(\mathbf{v}\cdot\mathbf{x})/\pi\pm\epsilon$ in time $\mathrm{poly}(d,1/\epsilon)$. Moreover, whenever $S$ is Gaussian our tests are guaranteed to pass (Theorem 2.4). The main challenge is that there are infinitely many potential vectors $v^{\prime}$ for which we want this property to hold, and brute-force search over all $v^{\prime}$ would take exponential time. We achieve this by discretizing $\mathbb{R}^{d}$ in buckets corresponding to $\mathbf{v}\cdot\mathbf{x}\in[i\epsilon,(i+1)\epsilon]$ for varying $i$ (Figure 2). We show that matching a constant number of moments for each bucket to the corresponding conditional Gaussian moments is sufficient for our purposes (Lemma 2.5), by using the notion of sandwiching polynomials for halfspaces to approximate the probability of disagreement in terms of bucket-wise low-degree polynomial functions (see Appendix B).
@@
Spectral tester and monotonicity under removal. The disagreement tester is only guaranteed to accept when the input $S$ is drawn i.i.d. from the standard Gaussian distribution. However, in our analysis it is important to have a tester that will accept even if given a set $S^{\prime}$ which is a subset of a Gaussian sample. We call this property monotonicity under removal and its importance is related to the fact that in the Massart noise model, the labels are not flipped independently of the corresponding features, but the probability of receiving a flipped label can adversarially depend on $\mathbf{x}$. To obtain a tester for the disagreement region that is monotone under removal (Theorem 2.9), we use an analysis similar to the one for the disagreement tester combined with ideas from the recent work by [GSSV24] on tolerant testable learning (see Appendix C). In particular, instead of matching low-degree moments, we bound the spectrum of the corresponding moment matrices (Algorithm 2).
@@ 1.3 Related Work
  • In time $d^{\tilde{O}(1/\epsilon^{2})}$ one can find a hypothesis with accuracy $\mathsf{opt}+\epsilon$ under Gaussian data distribution [KKMS08, DGJ+10]. See also [DKK+21] for a proper learning algorithm. An algorithm with a run-time $d^{O(1/\epsilon^{2-\Omega(1)})}$ (and let alone a polynomial run-time) is believed to be impossible due to statistical query lower bounds [GGK20, DKZ20, DKZ20], as well as recent cryptographic reductions from lattice problems [DKR23, Tie23]. This works utilize reductions to the continuous LWE problem [BRST21], shown in [GVV22] to be harder than the LWE problem widely used in lattice-based cryptography (a quantum reduction was given in [BRST21]).
  • A worse error bound of $O(\mathsf{opt})+\epsilon$ can be obtained in time $\mathrm{poly}(d/\epsilon)$ [ABL17]. Despite various refinements [Dan15, DKTZ20b, DKK+21], the improvement of the error bound to $\mathsf{opt}+\epsilon$ is precluded by the aforementioned hardness results.
@@
In order to obtain an error bound of $\mathsf{opt}+\epsilon$ in time $\mathrm{poly}(d/\epsilon)$ a large body of works focused on moving beyond worst-case models of label noise. In the Random Classification Noise (RCN) model [AL88] the labels are flipped independently with probability $\eta$. It was shown in [BFKV98] that in the RCN model halfspaces can be learned up to error $\mathsf{opt}+\epsilon$ in time $\mathrm{poly}(d/\epsilon)$. See also [Coh97, DKT21, DTK23, DDK+24].
@@
The Massart noise model, introduced in [MN06], is more general than the RCN model and allows the noise rate $\eta(\mathbf{x})$ to differ across different points $\mathbf{x}$ in space $\mathbb{R}^{d}$, as long as it is at most some rate $\eta_{0}$. First studied in [ABHU15], learning halfspaces up to error $\mathsf{opt}+\epsilon$ under Massart noise model has been the focus of a long line of work [ABHZ16, MV19, YZ17, ZLC17, DKTZ20a, ZL21, ZSA20, DKK+22].
@@
We would like to note that intermediate steps in the algorithm [DKK+22] work by finding what is referred in [DKK+22] as sum-of-squares certificates of optimality for certain halfspaces. We would like to emphasize that certificates in the sense of [DKK+22] have to be sound only assuming that the labels satisfy the Massart property. In contrast with this, certificates developed in this work satisfy soundness without making any assumptions on the label distribution (which is the central goal of this work).
@@
There has also been work on distribution-free learning under Massart noise [DGT19, CKMY20], which achieves an error bound of $\eta_{0}+\epsilon$, but as a result can lead to a much higher error than the information-theoretically optimal bound of $\mathsf{opt}+\epsilon$.
@@
Testable learning. The framework of testable learning was first introduced in [RV23] with a focus on developing algorithms in the agnostic learning setting that provide certificates of (approximate) optimality of the obtained hypotheses or detect that a distributional assumption does not hold. Many of the existing agnostic learning results have since been shown to have testable learning algorithms with matching run-times. This has been the case for agnostic learning algorithms with $\mathsf{opt}+\epsilon$ error guarantee [RV23, KSV24b, GSSV24, STW24], as well as $O(\mathsf{opt})+\epsilon$ error bounds [GGKS24, GKSV23, DKK+23, DKLZ24].
@@
We note that [GGKS24, GKSV23] also give testable learning algorithms in the setting where data labels are assumed to be Massart (and the algorithm needs to either output a hypothesis with error $\mathsf{opt}+\epsilon$ or detect that data distribution is e.g. not Gaussian). We emphasize that the results of [GGKS24, GKSV23] do not satisfy soundness when the user is not promised that data labels satisfy the Massart noise condition (which is the central goal of this work).
@@
TDS learning. The recently introduced TDS learning framework [KSV24b, KSV24a, CKK+24, GSSV24] considers a setting in which the learning algorithm is given a labeled training dataset and an unlabeled test dataset and aims to either (i) produce accurate labeling for the testing dataset (ii) detect that distribution shift has occurred and the test dataset is not produced from the same data distribution as the training dataset. Although conceptually similar, the work in TDS learning addresses a different assumption made in learning theory. Nevertheless, as we note, the spectral testing technique introduced in [GSSV24] is a crucial technical tool for our results in this work.

### v2

@@ Abstract
To model this question, we extend the recently proposed testable learning framework of Rubinfeld and Vasilyan [rubinfeld2022testing] and require a learner to run an associated test that satisfies the following two conditions: (1) whenever the test accepts, the learner outputs a classifier along with a certificate of optimality, and (2) the test must pass for any dataset drawn according to a specified modeling assumption on both the marginal distribution and the noise model. We then consider the problem of learning halfspaces over Gaussian marginals with Massart noise (where each label can be flipped with probability less than $1/2$ depending on the input features), and give a fully-polynomial time testable learning algorithm.
@@ 1 Introduction
Developing efficient algorithms for learning in the presence of noise is one of the most fundamental problems in machine learning with a long line of celebrated research. Assumptions on the noise model itself vary greatly. For example, the well-studied random classification noise model (RCN) assumes that the label corruption process is independent across examples, whereas malicious noise models allow a fraction of the (joint) data-generating distribution to be changed adversarially. Understanding the computational landscape of learning with respect to different noise models remains a challenging open problem, serving as the central focus of numerous works in the theory of supervised learning [blum1998polynomial, awasthi2015efficient, awasthi2016learning, yan2017revisiting, zhang2017hitting, mangoubi2019nonconvex, diakonikolas2020learning, diakonikolas2020learning, diakonikolas2022learning_general, diakonikolas2018learning, BSHOUTY2002255, diakonikolas2024near] and unsupervised learning [unsupervised_1, unsupervised_2, unsupervised_3, unsupervised_4, unsupervised_5, unsupervised_6, unsupervised_7, unsupervised_8, unsupervised_9, unsupervised_10, unsupervised_11, unsupervised_12].
@@
We use the recently introduced testable learning [rubinfeld2022testing] framework to model these questions. In this framework, a learner first runs a test on the training set. Whenever the test accepts, the learner outputs a classifier along with a proof that the classifier has near-optimal error. Furthermore, the test must accept with high probability whenever the training set is drawn from a distribution satisfying some specified set of modeling assumptions. If the test rejects, the learner recognizes that one of the modeling assumptions has failed and will therefore refrain from outputting a classifier. Here, our modeling assumptions will include both the structure of the noise model and the structure of the marginal distribution from which the data is generated.
More concretely, we will consider the problem of learning halfspaces under Gaussian marginals with respect to Massart noise, an extensively studied problem where an adversary flips binary labels independently with probability at most $1/2$ (the probability of flipping can vary across instances). The goal is to find a halfspace $\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})$ with near-optimal misclassification error rate $\mathsf{opt}+\epsilon$, where $\mathsf{opt}$ is the best misclassification error rate achievable by a halfspace. For this problem, a long line of work [awasthi2015efficient, awasthi2016learning, yan2017revisiting, zhang2017hitting, mangoubi2019nonconvex, diakonikolas2020learning] resulted in the algorithm of Diakonikolas et al. [diakonikolas2020learning] that runs in time $\mathrm{poly}(d/\epsilon)$ and achieves the optimal error rate. In contrast, the worst-case-noise version of this problem (i.e. agnostic learning or, equivalently, learning with adversarial label noise) is believed to require exponential time in the accuracy parameter, even with respect to Gaussian marginals [kalai2008agnostically, diakonikolas2021agnostic, goel2020statistical].
@@ 1.1 Our Results
Let $f:\mathbb{R}^{d}\to\{\pm 1\}$ be a concept, let $\eta:\mathbb{R}^{d}\to[0,1/2]$ and let $\mathcal{D}$ be a distribution over $\mathbb{R}^{d}$. The oracle $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{D},f,\eta}$ receives $m\in\mathbb{N}$ and returns $m$ i.i.d. examples of the form $(\mathbf{x},y)\in\mathbb{R}^{d}\times\{\pm 1\}$, where $\mathbf{x}\sim\mathcal{D}$ and $y=\xi\cdot f(\mathbf{x})$, with $\xi=1$ w.p. $1-\eta(\mathbf{x})$ and $\xi=-1$ w.p. $\eta(\mathbf{x})$. The quantity $\sup_{\mathbf{x}\in\mathbb{R}^{d}}\eta(\mathbf{x})\in[0,1/2]$ is called the noise rate.
@@
Learning Setting. Our results work in the following extension of testable learning [rubinfeld2022testing].
Definition 1.3 (Testable Learning, extension of Definition 4 in [rubinfeld2022testing]).
@@
The difference between Definition 1.3 and the definition of [rubinfeld2022testing] is that the completeness criterion does not only concern the marginal distribution on $\mathbb{R}^{d}$, but the joint distribution over $\mathbb{R}^{d}\times\{\pm 1\}$. The choice of the oracle class $\mathcal{O}$ encapsulates all of the modeling assumptions under which our algorithm should accept (both on the marginal distribution on $\mathbb{R}^{d}$, as well as on the labels). Note that the probability of success can be amplified through repetition (see [rubinfeld2022testing]), so it suffices to solve the problem for $\delta=1/3$. Our main results and their relation to prior work are summarized in Table 1.
@@
Massart, $\eta_{0}=1/2-c$ $\mathrm{poly}(d,1/\epsilon)$ [diakonikolas2020learning] $\mathrm{poly}(d,1/\epsilon)$ [Theorem 2.7]
@@
$d^{O(\log(1/\epsilon))}2^{\mathrm{poly}(1/\epsilon)}$ [diakonikolas2022learning_general]
$d^{\Omega(\log(1/\epsilon))}$ [diakonikolas2022learning_general]
$d^{\tilde{O}(1/\epsilon^{2})}$ [rubinfeld2022testing, gollakota2022moment]
@@
$d^{\tilde{O}(1/\epsilon^{2})}$ [kalai2008agnostically]
$d^{\Omega(1/\epsilon^{2})}$ [diakonikolas2021optimality]
$d^{\tilde{O}(1/\epsilon^{2})}$ [rubinfeld2022testing, gollakota2022moment]
@@
Upper Bound. In Theorem 2.7, we show that there is a polynomial-time tester-learner for the class $\mathcal{H}_{\mathsf{hs}}$ with respect to $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},\mathcal{H}_{\mathsf{hs}},\eta_{0}}$ for any $\eta_{0}\leq 1/2-c$, where $c$ is any positive constant. Moreover, whenever our algorithm accepts, it is guaranteed to output the optimal halfspace with respect to the input dataset $\bar{S}$, even if $\bar{S}$ is not generated from i.i.d. examples and can, therefore, be completely arbitrary. Given the upper bounds of Table 1 our algorithm can be used as a first step before applying the more powerful (but also more expensive) tester-learner of [rubinfeld2022testing, gollakota2022moment]. If our algorithm accepts, then we do not need to run the more expensive algorithm. In other words, our results highlight that testable learning can be used for algorithm selection for problems where different assumptions motivate different algorithmic approaches.
Lower Bounds. Our upper bound holds when the noise rate is bounded away below $1/2$. We show that this is necessary: in the high-noise regime ($\eta_{0}=1/2$), the best known lower bounds for learning under adversarial label noise also hold in the testable setting, with respect to random classification noise of rate $1/2$ (Definition 2.1), which is a special case of Massart noise. We give both cryptographic lower bounds (Theorem 3.4) assuming subexponential hardness on the problem of learning with errors (LWE), as well as statistical query lower bounds (Theorem 3.6). Our lower bounds are inherited from lower bounds from the literature of agnostic learning [diakonikolas2021optimality, tiegel2023hardness, diakonikolas2023near] (combined with 3.1). Our testable learning model highlights an underappreciated aspect of these agnostic learning lower bounds, namely, that the hard instances are in fact indistinguishable from completely random instances (i.e., random classification noise of rate $1/2$).
@@ 1.2 Our Techniques
The techniques we employ in this work are significantly more sophisticated than recently developed tools from testable learning. In fact, it is not even clear that techniques from testable learning should apply, as assumptions on the marginal distribution are quite different from assumptions on the noise model. Concretely, we depart from prior work in testable learning [gollakota2023efficient, gollakota2023tester] where the testers are designed to certify specific properties of a particular learning algorithm. Instead, here we obtain a “black-box” result that can take any learner that is guaranteed to output a near-optimal halfspace in the Massart setting and certify optimality properties of the learner’s output hypothesis. To do this, we decompose the error of the candidate output in terms of quantities for which we can provide certifiable bounds by developing appropriate testers (see Section 2 for more details on the decomposition). In particular, we provide a disagreement tester with significantly sharper guarantees compared to the one developed for standard testable learning [gollakota2023tester], and a spectral tester that combines and expands ideas from [gollakota2023tester] as well as recent work on tolerant testable learning [goel2024tolerant]. The main technical tool we develop to provide these improved guarantees is a notion of families of sandwiching approximators with respect to partitions of $\mathbb{R}^{d}$.
An outline of the proof of our main result (Theorem 2.7) is provided in Section 2. As a warm-up, in Section 2.1, we consider the special case of random classification noise, whose analysis is simpler. We complete the proof sketch of our main result in Section 2.2. In the following, we give an overview of the disagreement and the spectral testers.
Disagreement tester and sandwiching polynomials. Let $\mathbf{v}$ be a unit vector and $S$ be a dataset of size $\mathrm{poly}(d/\epsilon)$. If $S$ consists of Gaussian data-points, then for every unit vector $\mathbf{v}^{\prime}$ (w.h.p. over $S$)
$\operatorname*{\mathbb{P}}_{\mathbf{x}\in S}\left[\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})\right]=\measuredangle(\mathbf{v},\mathbf{v}^{\prime})/\pi\pm\epsilon.$
Suppose, given $S$ and $\mathbf{v}$, one would like to certify that this property approximately holds for every $\mathbf{v}^{\prime}$. The method of exhaustive search - i.e. checking this property for different candidate vectors $\mathbf{v}^{\prime}$ - can be shown to require at least $2^{\Omega(d)}$ time. Using a moment-based approach, we show how to improve this run-time exponentially. In particular, in time $\mathrm{poly}(d,1/\epsilon)$, we can certify that for all $\mathbf{v}^{\prime}\in\mathbb{S}^{d-1}$
$\operatorname*{\mathbb{P}}_{\mathbf{x}\in S}\left[\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})\right]=(1\pm 0.01)\measuredangle(\mathbf{v},\mathbf{v}^{\prime})/\pi\pm\epsilon.$ (1.1)
Moreover, whenever $S$ is Gaussian, our tests are guaranteed to pass (Theorem 2.4). Note that [gollakota2023efficient, gollakota2023tester, diakonikolas2023efficient] provided disagreement testers that certified one-sided bounds and suffered constant multiplicative error factors111i.e. certified that $\operatorname*{\mathbb{P}}_{\mathbf{x}\in S}\left[\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})\right]\leq O(\measuredangle(\mathbf{v},\mathbf{v}^{\prime}))+\epsilon$., while here our disagreement testers certify both upper and lower bounds on the disagreement probability, with a small and controllable multiplicative error factor.
As mentioned earlier, directly checking the disagreement for each candidate vector $\mathbf{v}^{\prime}$ in a Euclidean cover of the sphere $\mathbb{S}^{d-1}$ does not work, since their number is exponential to the dimension $d$. Instead, our tester discretizes $\mathbb{R}^{d}$ into buckets corresponding to $\mathbf{v}\cdot\mathbf{x}\in[i\epsilon,(i+1)\epsilon]$ for varying $i$ (Figure 2) and checks for any constant-degree monomial $m$ that
$\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[m(\mathbf{x})\cdot\mathbbm{1}_{i\epsilon\leq\mathbf{x}\cdot\mathbf{v}\leq(i+1)\epsilon}]\approx\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}[m(\mathbf{x})\cdot\mathbbm{1}_{i\epsilon\leq\mathbf{x}\cdot\mathbf{v}\leq(i+1)\epsilon}].$ (1.2)
We show that passing this test for constant-degree $m$ is sufficient for our purposes (Lemma 2.5). The previous work [gollakota2023efficient, gollakota2023tester, diakonikolas2023efficient] considered only tests involving monomials $m$ of degree at most $4$, and (as explained earlier) achieved bounds far weaker than Equation 1.1. A key ingredient to our improved testers is extending the notion of sandwiching polynomials of [gollakota2022moment] to much more general piecewise-polynomial functions (see Appendix B).
@@
Spectral tester and monotonicity under removal. The disagreement tester is only guaranteed to accept when the input $S$ is drawn i.i.d. from the standard Gaussian distribution. However, in our analysis it is important to have a tester that will accept even if given a set $S^{\prime}$ which is a subset of a Gaussian sample. We call this property monotonicity under removal and its importance is related to the fact that in the Massart noise model, the labels are not flipped independently of the corresponding features, but the probability of receiving a flipped label can adversarially depend on $\mathbf{x}$. Note that tester in Equation 1.2 is not monotone under removal.
To obtain a tester for the disagreement region that is monotone under removal (Theorem 2.9), we augment our Disagreement Tester using ideas from the recent work by [goel2024tolerant] on tolerant testable learning (see Appendix C). In particular, instead of checking Equation 1.2, our Spectral Tester checks that for every constant-degree polynomial $p$ we have
$\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[p(\mathbf{x})^{2}\cdot\mathbbm{1}_{i\epsilon\leq\mathbf{x}\cdot\mathbf{v}\leq(i+1)\epsilon}]\lesssim\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}[p(\mathbf{x})^{2}\cdot\mathbbm{1}_{i\epsilon\leq\mathbf{x}\cdot\mathbf{v}\leq(i+1)\epsilon}],$ (1.3)
which can be verified efficiently by computing the spectrum of an appropriate matrix. The main difference between our spectral tester and the one in [goel2024tolerant] is that ours partitions $\mathbb{R}^{d}$ into a number of strips and performs a test for each of them, while the one by [goel2024tolerant] runs the same test on the whole $\mathbb{R}^{d}$ iteratively, each time removing a number of points from the input set. See Algorithm 2 for the full algorithm description. As in the case of the Disagreement Tester, the analysis again leverages the method of piecewise-polynomial sandwiching functions introduced in this work.
@@ 1.3 Related Work
  • In time $d^{\tilde{O}(1/\epsilon^{2})}$ one can find a hypothesis with accuracy $\mathsf{opt}+\epsilon$ under Gaussian data distribution [kalai2008agnostically, diakonikolas2010bounded]. See also [diakonikolas2021agnostic] for a proper learning algorithm. An algorithm with a run-time $d^{O(1/\epsilon^{2-\Omega(1)})}$ (and let alone a polynomial run-time) is believed to be impossible due to statistical query lower bounds [goel2020statistical, diakonikolas2020near, diakonikolas2020near], as well as recent cryptographic reductions from lattice problems [diakonikolas2023near, tiegel2023hardness]. This works utilize reductions to the continuous LWE problem [bruna2021continuous], shown in [gupte2022continuous] to be harder than the LWE problem widely used in lattice-based cryptography (a quantum reduction was given in [bruna2021continuous]).
  • A worse error bound of $O(\mathsf{opt})+\epsilon$ can be obtained in time $\mathrm{poly}(d/\epsilon)$ [awasthi2017power]. Despite various refinements [daniely2015ptas, diakonikolas2020non, diakonikolas2021agnostic], the improvement of the error bound to $\mathsf{opt}+\epsilon$ is precluded by the aforementioned hardness results.
@@
In order to obtain an error bound of $\mathsf{opt}+\epsilon$ in time $\mathrm{poly}(d/\epsilon)$ a large body of works focused on moving beyond worst-case models of label noise. In the Random Classification Noise (RCN) model [angluin1988learning] the labels are flipped independently with probability $\eta$. It was shown in [blum1998polynomial] that in the RCN model halfspaces can be learned up to error $\mathsf{opt}+\epsilon$ in time $\mathrm{poly}(d/\epsilon)$. See also [cohen1997learning, diakonikolas2021forster, diakonikolas2023strongly, diakonikolas2024near].
@@
The Massart noise model, introduced in [massart2006risk], is more general than the RCN model and allows the noise rate $\eta(\mathbf{x})$ to differ across different points $\mathbf{x}$ in space $\mathbb{R}^{d}$, as long as it is at most some rate $\eta_{0}$. First studied in [awasthi2015efficient], learning halfspaces up to error $\mathsf{opt}+\epsilon$ under Massart noise model has been the focus of a long line of work [awasthi2016learning, mangoubi2019nonconvex, yan2017revisiting, zhang2017hitting, diakonikolas2020learning, zhang2021improved, zhang2020efficient, diakonikolas2022learning_general].
@@
We would like to note that intermediate steps in the algorithm [diakonikolas2022learning_general] work by finding what is referred in [diakonikolas2022learning_general] as sum-of-squares certificates of optimality for certain halfspaces. We would like to emphasize that certificates in the sense of [diakonikolas2022learning_general] have to be sound only assuming that the labels satisfy the Massart property. In contrast with this, certificates developed in this work satisfy soundness without making any assumptions on the label distribution (which is the central goal of this work).
@@
There has also been work on distribution-free learning under Massart noise [diakonikolas2019distribution, chen2020classification], which achieves an error bound of $\eta_{0}+\epsilon$, but as a result can lead to a much higher error than the information-theoretically optimal bound of $\mathsf{opt}+\epsilon$.
@@
Testable learning. The framework of testable learning was introduced in [rubinfeld2022testing] with a focus on developing algorithms in the agnostic learning setting that provide certificates of (approximate) optimality of the obtained hypotheses or detect that a distributional assumption does not hold. Many of the existing agnostic learning results have since been shown to have testable learning algorithms with matching run-times. This has been the case for agnostic learning algorithms with $\mathsf{opt}+\epsilon$ error guarantee [rubinfeld2022testing, klivans2023testable, goel2024tolerant, slot2024testably], as well as $O(\mathsf{opt})+\epsilon$ error bounds [gollakota2024agnostically, gollakota2023tester, diakonikolas2023efficient, diakonikolas2024testable].
@@
We note that [gollakota2024agnostically, gollakota2023tester] also give testable learning algorithms in the setting where data labels are assumed to be Massart (and the algorithm needs to either output a hypothesis with error $\mathsf{opt}+\epsilon$ or detect that data distribution is e.g. not Gaussian). We emphasize that the results of [gollakota2024agnostically, gollakota2023tester] do not satisfy soundness when the user is not promised that data labels satisfy the Massart noise condition (which is the central goal of this work).
@@
Testable Learning with Distribution Shift (TDS learning). The recently introduced TDS framework [klivans2023testable, klivans2024learning, chandrasekaran2024efficient, goel2024tolerant] considers a setting in which the learning algorithm is given a labeled training dataset and an unlabeled test dataset and aims to either (i) produce accurate labeling for the testing dataset (ii) detect that distribution shift has occurred and the test dataset is not produced from the same data distribution as the training dataset. Although conceptually similar, the work in TDS learning addresses a different assumption made in learning theory. Nevertheless, as we note, the spectral testing technique introduced in [goel2024tolerant] is a crucial technical tool for our results in this work.